OpenStudy (anonymous):

how do i solve 5 over 7 = x over 98?

OpenStudy (saifoo.khan):

cross multiply.

OpenStudy (anonymous):

cross multiply 7x=490 x=70

OpenStudy (saifoo.khan):

98 x 5 = 7 x x

OpenStudy (anonymous):

7x = 490 ?

OpenStudy (saifoo.khan):

490/7

OpenStudy (saifoo.khan):

x = 70

OpenStudy (anonymous):

I WAS FIRST

OpenStudy (saifoo.khan):

lol

OpenStudy (anonymous):

so x over 5 = 12 over 30 would be 60 = 30x then 60/30=2 = 30/30 =x... so x=2?

myininaya (myininaya):

yes